A 5.0-µC point charge is placed at the 0.00 cm mark of a meter stick and a -4.0-µC point charge is placed at the 50 cm mark. At what point on a line joining the two charges is the electric field due to these charges equal to zero?

Draw a sketch of the charge distribution. • Establish a coordinate system and show the locations of the charges. • Identify the point P at which you want to calculate the electric field. • Draw the electric field of each charge at point P. • Use symmetry to determine if any components of E net are zero.

For each charge, determine its distance from P and the angle of Ei • Determine the field strength of each charge. • Write each vector in component form. • Sum the vector components to determine E net . • Check result for correct units and that it is reasonable.
Answer #1

+y akisc î) + 5yc -4uc Р rî) xeo 250 E En +2 ares Soom r2 m. = 50+82)com Electric field due to Q1 kQ E Electric field due to5 4 2 (50+r2) (2) 50+r2 | || T2 2 100 + 2r2 = /5r2 100 2 V5-2 r2 = 423.6cm ri = 50 + 423.6 = 473.6cm